OpenStudy (anonymous):

Find the larger root for 7x^2− 2 = 2

jhonyy9 (jhonyy9):

7x^2 =4 x^2 =4/7 x_1,2=+/- sqrt(4/7) = +/- (2sqrt7)/7

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

i dont get it

jhonyy9 (jhonyy9):

first you need to assum to both sides 2 after divide both sides by 7

jhonyy9 (jhonyy9):

- so than will get x^2 = 4/7 x_1,2 =+/- sqrt(4/7) = +/- (srt4)/sqrt7) =+/- 2/(sqrt7) = +/- (2sqrt7)/7